WARNING: This web is not used anymore. Please use PDBService.MinuteS3July08 instead!
 

Minutes 3 July 08

  • Phone:
    • Gridka: Andreas, Doris, Camile
    • NDGF: Olli
    • IN2P3: Jerome
    • PIC: Luis

  • CERN:
    • 3D: Dirk, Maria, Eva
    • ATLAS: Gancho, Florbela

  • Today cannot join - reports sent by email:
    • BNL: Prepare space and account for TAGS tests. Reconfigure memory parameters current SGA 4G, db_keep_cache_size to 1200M node 2 (this parameter is not needed anymore – Florbela). Stream pool was not modified on this node.
    • SARA: No problems to report. We didn't upgrade Oracle yet but we made all necessary preparations. We expect to upgrade somewhere in the next two weeks.
    • Taiwan: We have upgraded 3D Oracle RAC and the agent to 10.2.0.4 this week.

  • Sites updates:
    • Gridka (Andreas): Upgrade to 10.2.0.4 planned to July 9th. Project leader has changed at GridKa.
    • IN2P3: 2 new RAC clusters dedicated to LHCb and ATLAS and AMI databases: quadcore boxes and Linux 64bit. 15G RAM.
      • Hw will be ready mid next week.
      • Databases migration – using transportable tablespaces for ATLAS, datapump import and export for LHCb and LFC.
      • AMI very critical, requesting high availability. Streams to be implemented.
    • PIC (Luis): Upgrade to 10.2.0.4 planned next 8th July. Working with Florbela for TAGs tests. Preparing tablespace.
    • NDGF (Olli): Database upgrade to 10.2.0.4 on 2nd July.
      • Parameter aq_tm_processes parameter setup during upgrade using dbua assistant
      • This parameter must not be set explicitly. Please check and remove it from the spfile/pfile
      • Review metalink Note:428441.1https://twiki.cern.ch/twiki/bin/view/PSSGroup/StreamsRecommendations
      • Olli will be on holidays for the next 4 weeks.
    • CERN:

  • TAGS:
    • Florbela sent document with the upload rate.
    • Index rebuilding – On INT8R data can be inserted with the indexes enabled. This is not the case with Triumf, indexes must be disabled and re-building, probably related with the SGA size.
    • Tests to be done on BNL and PIC during the next weeks.
    • DESY is preparing the hw.

  • PVSS:
    • Working on the first account to be replicated in production. Some problems with update statements, working on it.
    • Maria asked for PVSS problems running in 10.2.0.4. Atlas has not experience any problem since the upgrade.

Meeting in 2 weeks: 17th July

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-07-09 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback